### 13 Dec 2023 v1

# Fracton models from product codes

Abstract
We explore a deep connection between fracton order and product codes. In particular, we propose and analyze conditions on classical seed codes which lead to fracton order in the resulting quantum product codes. Depending on the properties of the input codes, product codes can realize either Type-I or Type-II fracton models, in both nonlocal and local constructions. For the nonlocal case, we show that a recently proposed model of lineons on an irregular graph can be obtained as a hypergraph product code. Interestingly, constrained mobility in this model arises only from glassiness associated with the graph. For the local case, we introduce a novel type of classical LDPC code defined on a planar aperiodic tiling. By considering the specific example of the pinwheel tiling, we demonstrate the systematic construction of local Type-I and Type-II fracton models as product codes. Our work establishes product codes as a natural setting for exploring fracton order.
Recent years have seen ideas and techniques from quantum information drive new developments in many areas of physics. Among these are new insights around an old problem in statistical mechanics, namely, the physics of glassiness. Exotic models with glassy phenomenology typify both the emerging field of fracton topological order [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] as well as breakthrough developments of asymptotically good quantum low-density parity check (LDPC) codes implemented in non-local stabilizer Hamiltonians [13, 14, 15].
Figure 1: (a) The schematic depicts set inclusions for models representable by either product codes or the polynomial formalism [16, *Haah_Algebraic2016]. Explicit product code constructions of certain models are provided in the supplementary material [18]. (b,c) Classical Tanner graphs are shown, with bits represented by blue circles and checks by red squares. The HGP is represented graphically: in (b), a surface code is found as the HGP of two repetition codes, and in (c), the anisotropic $\mathbb{Z}_{2}$ Laplacian model [19] is obtained from a repetition code and a Laplacian code.
At first glance, these two topics seem rather distinct. On the one hand, fracton phases of matter constitute a fascinating counterexample to the lore of solid state physics by reproducing glassy dynamics within local models exhibiting translation invariance [1]. Features distinguishing such phases from both symmetry breaking and conventional topological order include: (i) ground-state degeneracy in finite systems that depends nontrivially on the volume, and (ii) low-energy excitations with constrained or no mobility. Instances with constrained mobility are referred to as Type-I fracton models, and those with no mobility as Type-II fracton models [2, 3, 4, 5].
On the other hand, quantum LDPC codes represent a promising type of quantum error correction that can be defined via a stabilizer Hamiltonian $\mathcal{H}=\sum_{j}\Pi_{j}$, where each term $\Pi_{j}$ has bounded norm and acts on a constant number of qubits, and each qubit participates in only a constant number of terms. A systematic approach to generating such codes is the family of product constructions, which take as input multiple classical LDPC codes and output a quantum LDPC code [20, 21, 13]. The canonical example is the hypergraph product (HGP), a prescription for generating a CSS quantum code from a pair of classical codes [22]. Crucially, product codes allow analytic calculation of the resulting code parameters including the encoding rate (the number of logical qubits per physical qubit) and the code distance (the minimal number of errors required to connect logical states).
Despite their different contexts, these topics are connected by certain individual results: for instance, Haah’s code, a Type-II fracton model, can be obtained via a generalization of the HGP known as the lifted product [21, 14]. One can also use classical codes to construct quantum “fractal spin liquids” [3], based on the polynomial formalism for translation invariant codes [16, *Haah_Algebraic2016] (Fig. 1).
More broadly, an interesting observation suggests a fundamental connection between fracton order and HGP codes. To see this, we note that, for gapped phases of matter, the logarithm of the ground-state degeneracy determines the number of qubits that could in principle be encoded in the ground space. In conventional phases, this effective qubit number is independent of system size. This scaling is evidently incompatible with the number of logical qubits in typical HGP codes, which is polynomial in the system size. By contrast, such polynomial scaling is reminiscent of precisely the nontrivial dependence of the ground-state degeneracy on system size in fracton phases. Despite this tantalizing connection, a sharp understanding of the relationship between fracton order and product codes remains lacking.
In this Letter, we provide such a framework, and demonstrate that product codes provide a natural setting for discovering new fracton phases of matter (Fig. 1). Our main results are threefold. First, we extend the definition of fracton order to stabilizer models without translation invariance or geometric locality. Based on this definition, we propose a set of criteria on the input classical codes of a product construction. When satisfied, the resulting quantum LDPC code realizes a nonlocal fracton phase. Second, we show that a recent model, purported to exhibit Type-I fracton order on an irregular graph, can be obtained as a hypergraph product code. Interestingly, while this particular case does not, in fact, constitute a fracton model, a generalized construction (utilizing a typical LDPC input code) does.
Finally, we consider geometrically local product constructions, introducing a new family of classical LDPC codes defined on aperiodic tilings of the plane. Specifically, we consider a local two-dimensional code on the pinwheel tiling (Fig. 3), which we term the pinwheel code. We show that the HGP of the pinwheel code with a repetition code produces a local Type-I fracton model in three dimensions, and that the HGP of two pinwheel codes produces a local Type-II fracton model in four dimensions. These results open the door to the systematic discovery of new local fracton models using product codes. Somewhat intriguingly, the resulting models also exhibit evidence of a feature known as confinement (see discussions below), a favorable property of certain error correcting codes that is violated in most known fracton models [23, 24].
Hypergraph product codes.—A classical LDPC code on $n$ bits with $m$ terms is represented by a $\kappa$-sparse $m\times n$ parity check matrix $H$ with elements in $\mathbb{F}_{2}=\{0,1\}$. It is characterized by a triplet $[n,k,d]$, where $k$ is the number of encoded logical bits and $d$ is the code distance [25]. The number of bits participating in each check is bounded above by a constant $\kappa_{c}$, and the number of checks acting on any given bit is bounded by $\kappa_{b}$; then $\kappa=\max(\kappa_{b},\kappa_{c})=O(1)$. Typical instances of LDPC codes are obtained by randomly sampling parity check matrices $H$ [18].
Similarly, a CSS quantum code is represented by two parity check matrices $H_{X}$ and $H_{Z}$ specifying the action of the $X$ and $Z$ stabilizers, respectively [26, 27]. Both classical and quantum parity check matrices are maps taking a configuration of errors, represented by a $n$-dimensional vector over $\mathbb{F}_{2}$, to a syndrome or excitation, represented by an $m$-dimensional vector over $\mathbb{F}_{2}$.
The hypergraph product (HGP) acts on a pair of classical “seed” codes, producing a CSS quantum code which is LDPC if the inputs are LDPC [22]. The quantum code resulting from the HGP of classical seed codes $H_{1}$ (with $n_{1}$ bits and $m_{1}$ checks) and $H_{2}$ (with $n_{2}$ bits and $m_{2}$ checks) is given by:
$\begin{gathered}H_{X}=\begin{bmatrix}H_{1}\otimes\mathbb{I}_{n_{2}}&\mathbb{I}_{m_{1}}\otimes H_{2}^{\top}\end{bmatrix},\\ H_{Z}=\begin{bmatrix}\mathbb{I}_{n_{1}}\otimes H_{2}&H_{1}^{\top}\otimes\mathbb{I}_{m_{2}}\end{bmatrix},\end{gathered}$ (1)
with $\mathbb{I}$ being the identity matrix. Classical seed codes with parameters $[n_{1},k_{1},d_{1}]$ and $[n_{2},k_{2},d_{2}]$ produce a quantum code with parameters $[[n_{q},k_{q},d_{q}]]$, where [22]
$\begin{gathered}n_{q}=n_{1}n_{2}+m_{1}m_{2}~{},\\ k_{q}=k_{1}k_{2}+k_{1}^{\top}k_{2}^{\top}~{},\\ k_{X}^{\top}=k_{1}^{\top}k_{2}~{},~{}k_{Z}^{\top}=k_{1}k_{2}^{\top}~{},\\ d_{q}=\min(d_{1},d_{2},d_{1}^{\top},d_{2}^{\top})~{}.\end{gathered}$ (2)
Here, $k_{i}^{\top}$ and $d_{i}^{\top}$ denote the parameters of the transposed (or dual) code, with parity check matrix $H_{i}^{\top}$. The number of superselection sectors in the quantum code is given by $2^{k_{X}^{\top}+k_{Z}^{\top}}$ [18]. The prototypical example of a hypergraph product is the toric code, obtained from two cyclic repetition codes having $k_{1}=k_{1}^{\top}=k_{2}=k_{2}^{\top}=1$ and $d_{1}=n_{1}$, $d_{2}=n_{2}$ [22].
As shown in Fig. 1(b), the HGP admits an intuitive graphical representation based on the Tanner graph; a bipartite Tanner graph describes any classical code by representing both checks and bits as vertices, with edges denoting the connectivity of the terms in $H$ [28].
Fracton order in nonlocal product codes.—Product constructions of stabilizer codes generically do not feature spatial symmetry or geometric locality; to connect to conventional fracton models [11, 12] and the local codes that we discuss later, we begin by providing a concrete definition of nonlocal fracton order. In particular, to ensure a well-defined thermodynamic limit, we sample from a random ensemble specifying a family of codes with connected Tanner graphs as $n_{q}\to\infty$  [^1] We have in mind codes which are far away from geometric locality, rather than hybrid geometries [55, 56].. Such a family is said to host fractons if: the logarithm of the ground state degeneracy scales polynomially in the system size  [^2] Note that this scaling is distinct from that of local fractons, which merely follows a polynomial envelope.; and excitations are pointlike  [^3] Finite-dimensional local fracton models with immobile loop excitations are known [57]. However such a case is not expected for the nonlocal models we consider. and have constrained mobility due to superselection sectors which also scale polynomially in the system size. An additional requirement in the case of nonlocal Type-II fractons is “confinement” [24], a natural consequence of nonlocality that implies the absence of string-like operators  [^4] The definition of confinement here is weaker than its usage elsewhere [23], as the energy cost is allowed to scale as any increasing function of the error weight, rather than strictly linearly.. The essence of confinement is that for sparse error patterns, additional errors always increase the number of syndromes.
Product constructions provide the opportunity to reason about quantum codes through investigation of their classical seeds. We propose three criteria for the classical seed codes that lead to nonlocal fracton order in the resulting HGP code: (i) rank deficiency, (ii) confinement, and (iii) isolability.
Rank deficiency refers to a property of the parity check matrix $H$. Since codewords are zero eigenvectors, a rank-deficient $H$ implies a large code space, with $k=n-\operatorname{\mathrm{rank}}(H)$. A rank-deficient seed code gives rise to superselection sectors in the resulting quantum code. A superselection sector is an equivalence class of syndromes under local errors; as the number of sectors grows, a diminishing number of excitation configurations may be dynamically connected by any physical process. This constitutes a strong mechanism for immobility, and is central to both local and nonlocal fracton models [33].
Intuitively, such an effect can be compared to the role of superselection sectors in, for example, the toric code, where they are labeled by elements of the anyon theory [34, 35]. Both cases identify dynamically disconnected excitation patterns; the difference is that topologically ordered phases like the toric code support only finitely many superselection sectors, whereas in an HGP code based on a rank-deficient seed code, the number of superselection sectors grows with the system size.
Another source of immobility is confinement, a property related to glassiness which is studied in both classical and quantum codes [23, 24]. A typical classical LDPC code is linearly confining with high probability, as nonlocal checks lead individual errors to produce roughly $\kappa_{b}$ excitations. Except in special cases [3], if both seed codes are confining, then so is the resulting quantum code.
An example of a classical seed code satisfying the prior conditions is the nonlocal Ising model; however the HGP of two such codes is not a fracton model. This is a consequence of supporting only excitations which are extended objects [23], however, and is special within the space of nonlocal codes. In order to exclude these cases, we impose an isolation condition requiring an individual excitation to be point-like; that is, it can be separated from other excitations with high probability.
Isolability is a natural feature of nonlocal codes, as a simple probabilistic argument demonstrates [18]. This can be seen as a consequence of the unlikelihood of an “Ising subgraph”: a set of bits on the Tanner graph whose neighboring checks are all two-valent. Such an Ising subgraph contains a set of local symmetries constraining excitations to be extended domain walls, and thus not isolable  [^5] In particular, this refers to a set of classical meta-checks, one for each independent cycle in the subgraph.. Crucially, typical LDPC codes with $\kappa_{c}>2$ are isolable, as the likelihood of large Ising subgraphs is exponentially suppressed. As was the case with confinement, a HGP code inherits isolability only if this condition is satisfied in both seed codes.
Nonlocal Type-I fractons in HGP codes.—We consider hypergraph product codes, in which one classical seed is a cyclic repetition code and the other is either: case (i), a Laplacian code; or case (ii), a typical LDPC code (see [18] for details of each). This is motivated by a recently proposed stabilizer model for Type-I fracton order called the “anisotropic $\mathbb{Z}_{2}$ Laplacian model” [37]. While this model was originally derived from a series of results on dipole-conserving tensor gauge theories [38, 39, 19], it turns out to be precisely described by case (i) above [Fig. 1(c)] [18].
Figure 2: (a) Rank deficiency is exhibited for typical classical LDPC codes, from the linear scaling of $k$ with $n$; by contrast, the number of logical bits in a Laplacian code is on average bounded. (b) Confinement data is shown for both typical LDPC codes (blue) and Laplacian codes (red), on systems of $n=300$ and $500$ bits. Pictured is the minimum syndrome weight observed for randomly-sampled error configurations. Both constructions display evidence of confinement, as expected for nonlocal codes.
The parity check matrix of the classical Laplacian code used in case (i) is given by the reduction mod 2 of the Laplacian operator of a graph [40]. To study its rank deficiency, we find the number of encoded logical bits $k$ as a function of system size $n$. For each system size, we average over $\sim 10^{3}$ random graphs, showing the results in Fig. 2(a). It is immediately evident that the Laplacian code on a general sparse graph does not exhibit evidence of rank deficiency. As a consequence, the resulting quantum code does not support fracton order.
Figure 3: (a) The pinwheel tiling substitution rule, taking one triangle to five, is shown. The graphs considered here are generated from two base tiles. (b,c) Graphs $G_{N}$ at generations $N=3,4$ are shown for the pinwheel tiling on a rectangular region. Each vertex hosts both a bit and a check, except for a fraction of boundary sites, whose checks are removed. The lowest-weight codeword is indicated in red for each. (d) The scaling of the number of logical bits $k\sim\sqrt{n}$ as well as the distance $d\sim n$ are shown for the pinwheel code, for several values $p=7,11,15$. Distances shown are exact, with the exception of the largest value having $k=65$, for which an upper bound is shown. (e) The minimum syndrome weight over sampled errors at $N=5$ provides numerical evidence for confinement in this local model. For low-weight errors, all error configurations are considered within balls of fixed radius centered on randomly sampled vertices. For high-weight errors, local truncations of logical operators illustrate the naive square-root upper bound. Though not exhaustive, such sampling can diagnose lack of confinement in practice [18].
This is in apparent contrast with results on the square lattice [37], but can be understood from the perspective of the seed code. Indeed, typical LDPC codes used in case (ii) are chosen with an imbalance between bits and checks leading to $k=O(n)$ rank deficiency, as shown in Fig. 2(a). Because the parity check matrix of the Laplacian code is square, it is very likely to be nearly full rank. The square lattice is thus a special case, and exact results can in fact illuminate how translation invariance leads to rank deficiency in the Laplacian code [18].
In contrast, we find that confinement is a feature of both typical LDPC and Laplacian codes. We compute the minimum density $|s|/m$ of excitations induced by $\sim 10^{3}$ randomly sampled error configurations at each sparsity $|e|/n$. The monotonic growth illustrated in Fig. 2(b) suggests that both typical LDPC and Laplacian codes are confining. Finally, we observe that both typical LDPC and Laplacian codes are isolable due to the suppression of Ising subgraphs by taking $\kappa>2$ [18].
Thus, we conclude that typical LDPC codes are rank-deficient, confining, and isolable, whereas Laplacian codes on general graphs are merely confining and isolable. Because the cyclic repetition code is also isolable, our case (ii) HGP code indeed realizes Type-I fractons. However, rather than fractons, the case (i) HGP construction instead hosts partially confined point-like excitations in a phase akin to nonlocal topological order.
Local fracton models from aperiodic codes.— Though the stabilizer models discussed so far are only $\kappa$-local, it is natural to ask whether geometrically local HGP codes realize local fracton order. Whereas one well-known route to locality is to define a Tanner graph on a negatively curved manifold, which can embed expanders [41], we instead take a Euclidean approach by placing both bits and checks on the vertices of a two-dimensional aperiodic tiling.
As a specific example, we consider the pinwheel tiling, which has no exact spatial symmetries [42, 43, 44, 45]. This tiling is specified by a substitution rule [Fig. 3(a)] generating a family of graphs $\{G_{N}\}$, $N=1,2,\ldots$, whose volume scales as $n\sim e^{N}$ [see Fig. 3(b,c)]. We define parity check matrices $H_{N}$ based on the graph Laplacian $L_{N}$ at generation $N$ [40]. We first compute $\tilde{H}_{N}=(L_{N}-\mathbb{I})\mod 2$, where $\mathbb{I}$ is the $n\times n$ identity matrix. Then $H_{N}$ is obtained via a “boundary depletion” of checks, truncating the column space of $\tilde{H}_{N}$ by removing an evenly spaced fraction $1/p$ of checks on the boundary vertices of $G_{N}$. We dub this the pinwheel code, within a more general class of aperiodic codes.
At issue are thus the rank deficiency and confinement of the pinwheel code; isolability is guaranteed, as the connectivity of bits and checks ranges between three and nine for all generations. As shown in Fig. 3(d), this family exhibits $k\approx\sqrt{n}/p$ and is thus rank deficient, with the parameter $p$ tuning the number of logical bits  [^6] We observe that for small values of $p$ short logical operators attached to the boundary can appear, thus we take large enough values to suppress these.. Interestingly, although the depletion of checks relative to bits is a boundary effect, this does not imply that logical bits are localized on the boundary [18]. Indeed, we find that this is not the case, measuring a code distance that scales linearly with $n$, as shown in Fig. 3(d). The pinwheel code introduced here thus saturates the classical BPT bound $k\sqrt{d}=O(n)$ [47].
In contrast to typical LDPC codes, a $D$-dimensional local code cannot exhibit linear confinement. An upper bound is realized by the Ising model, wherein domain walls have polynomial energy cost with exponent $\frac{D-1}{D}$; more generally, in any local code, a codeword can be truncated within a $D$-ball at the cost of only a $(D-1)$-dimensional boundary syndrome. Such scaling is exhibited in the limit of large error weights in Fig. 3(e), and indeed follows a power law with exponent $1/2$. Though sampling is not guaranteed to capture the tails of the distribution, it is sufficiently thorough to distinguish the present cases from the Laplacian code on the square lattice, which, though rank deficient, is not confining [18]. We thus conjecture that the pinwheel code is confining, and saturates the upper bound.
Having demonstrated rank deficiency, confinement, and isolability, we now consider producing local fracton models via the hypergraph product of pinwheel codes. First, taking as inputs a pinwheel code and a cyclic repetition code yields a local Type-I fracton model in three dimensions. This code exhibits translation invariance only in the out-of-plane direction (i.e., perpendicular to the tiling). Second, taking the seed codes to be two instances of pinwheel codes yields a local Type-II fracton model in four dimensions, with parameters $k_{q}\sim n_{q}^{1/4}$, $d_{q}\sim\sqrt{n_{q}}$  [^7] While we have studied a particular rectangular patch under the substitution rules, more general regions are possible, and due to the lack of spatial symmetry non-identical codes will not have large coincident subgraphs.. We therefore largely capture the physics associated with the typical LDPC code [case (ii)] without having to appeal to nonlocal interactions or disorder. We conjecture that, similar to the four-dimensional Ising model, the Type-II pinwheel stabilizer model is confining without disorder, but, crucially, achieves this for point-like rather than extended excitations.
Discussion and conclusion.— From the perspective of the present work, the study of quantum glassiness leading to fracton order can be understood as an exchange of confinement for superselection, each of which provides a complementary approach to constraining mobility [33]. Here, we consider neither to be fundamentally classical or quantum, and instead characterize both as aspects of quantum product codes.
Looking forward, our work opens the door to a number of intriguing directions. First, more sophisticated product codes include the lifted and balanced products, and achieve a dimensional reduction by making use of lattice symmetries of the seed codes [20, 21, 13]. It is not immediately evident how to extend such an operation to aperiodic codes; however, this could in principle reduce a code similar to our local Type-II model to three dimensions, which could be compared to other recent constructions [49]. Second, it would be interesting to characterize the conditions on aperiodic tilings that lead to local Type-II fractons, and whether such tilings all flow to the same fixed point under entanglement renormalization [50, 51, 52, 53]. One possible avenue is an understanding in terms of condensation of certain topological lattice defects [54]. Finally, it would be useful to quantitatively investigate the spin-glass contribution to immobility and the energy barrier in product code fracton models.
Acknowledgements.—We gratefully acknowledge the insight of and discussions with Hong-Ye Hu, Zhu-Xi Luo, Shu-Heng Shao, Thomas Schuster, and Carolyn Zhang. We are particularly indebted to Helia Kamal for discussions leading to the origination of this project. This work was supported in part by multiple grants from the U.S. Department of Energy, including the Accelerated Research in Quantum Computing (ARQC) program, the Quantum Information Science Enabled Discovery (QuantISED) for High Energy Physics (KA2401032) and through the GeoFlow Grant No. de-sc0019380. B. R. acknowledges support from the Harvard Quantum Initiative Postdoctoral Fellowship in Science Engineering. N. Y. Y. acknowledges support from a Simons Investigator award.